Find the solution of the following symmetrical simultaneous differential equation dx/1 = -dy/1 = dz/1


1
Expert's answer
2022-02-01T16:21:19-0500

"dx\/1 = -dy\/1 = dz\/1"

"dx = -dy = dz"

"d(x+y)=d(z+y)=0"

Therefore, the functions "x+y" and "z+y" are constant along any integral curve and form a complete system of integral invariants. Therefore, the general solution can be written as

"\Phi(x+y,z+y)=0", where "\Phi" is an arbitrary smooth function of two variables.